### v1

# Unsupervised Meta-Learning via In-Context Learning

Abstract
Unsupervised meta-learning aims to learn feature representations from unsupervised datasets that can transfer to downstream tasks with limited labeled data. In this paper, we propose a novel approach to unsupervised meta-learning that leverages the generalization abilities of in-context learning observed in transformer architectures. Our method reframes meta-learning as a sequence modeling problem, enabling the transformer encoder to learn task context from support images and utilize it to predict query images. At the core of our approach lies the creation of diverse tasks generated using a combination of data augmentations and a mixing strategy that challenges the model during training while fostering generalization to unseen tasks at test time. Experimental results on benchmark datasets, including miniImageNet, CIFAR-fs, CUB, and Aircraft, showcase the superiority of our approach over existing unsupervised meta-learning baselines, establishing it as the new state-of-the-art in the field. Remarkably, our method achieves competitive results with supervised and self-supervised approaches, underscoring the efficacy of the model in leveraging generalization over memorization.
**footnotetext: Equal contribution
1 Introduction
Meta-learning, or learning-to-learn, empowers learning systems with the ability to accumulate knowledge from multiple tasks, enabling rapid adaptation and generalization to new tasks [48, 45]. Traditional meta-learning approaches typically rely on labeled data to construct tasks during meta-training. However, collecting large labeled datasets in real-world applications is challenging and often impractical. Addressing this issue, unsupervised meta-learning (UML) approaches aim to leverage unlabeled data to learn transferable feature representations, enabling adaptation to new tasks with limited labeled data [48].
Various approaches have been proposed to address the UML problem [18, 19, 20, 23, 24, 25]. Despite its potential, UML faces several challenges. Existing UML methods often rely on simple data augmentations to construct the training tasks, while following the standard meta-learning task sampling pipeline for the evaluation phase. This results in a significant difference between training and testing tasks, constraining the model’s ability to generalize to unseen tasks during inference and requiring additional fine-tuning on the test domain. Furthermore, existing UML approaches typically assume that the training and test datasets belong to the same domain. In this work, we loosen this assumption resulting in a more challenging setting that necessitates better model generalization compared to usual meta-learning applications. We refer to this as the cross-domain scenario.
In this paper, we propose a novel approach to UML that addresses these challenges by leveraging in-context learning within a transformer architecture [11, 31]. In-context learning [11] allows the model to use the context provided by a sequence of input-output pairs to make predictions about new inputs. Inspired by recent advancements in large language models (LLMs) [52, 4, 29], we formulate meta-learning as a sequence modeling problem, where a task is seen as a sequence of support images and an unknown query image. The support set is treated as the context utilized by the model to predict the class of the query image. We call our approach CAMeLU, which stands for Context-Aware Meta-Learning in Unsupervised scenarios. Central to our approach is a novel task creation mechanism that enables the generation of a large number of different tasks from an unlabeled dataset. Drawing inspiration from the natural decision process of learning by analogy [53], we construct tasks that closely resemble the structure of those encountered during inference. Specifically, we use a combination of different data augmentation techniques based on basic image manipulations [39] for generating the samples in the support set. Conversely, a strategy similar to mixup [57] is employed to generate query images by combining a support element and an image randomly sampled from the training dataset. This process ensures that the query contains sufficient information from the support image to be classified as the latter, while introducing diversity by blending them. Consequently, query images appear distinct from their corresponding support images while still belonging to the same class, better mimicking the tasks seen at test time and hence enhancing generalization. Following task creation, support and query images are encoded using a fixed pre-trained feature extractor. The resulting latent representations are aggregated into a sequence and passed as input to a transformer encoder along with their label encodings. The transformer encoder learns to extract contextual information from support images and predict the query image in a single pass, eliminating the need for the fine-tuning step during inference. An overview of our approach is visualized in Figure 1.
Figure 1: Visualization of CAMeLU (example with 3-way 5-shot tasks). On the left side is the task creation mechanism, where $N$ samples are sampled from an unlabeled dataset $\mathcal{D}_{train}$. Each sample $x_{n}$ is augmented $K$ times to obtain $x_{n,k}$, and the same pseudo-label $n\in[1,N]$ is assigned to all data generated from the sample $x_{n}$. A strategy inspired by mixup [57] is utilized for generating the query set. On the right side, the so-created task is fed to the transformer encoder for predicting the query input. Inspired by CAML [12], we fed the transformer encoder with demonstrations built upon a concatenation of the features extracted from a fixed pre-trained feature extractor and a learned class encoder. The symbol * represents the unknown query label that the transformer encoder aims to predict.
Throughout extensive experiments we demonstrate the effectiveness of the proposed approach to generalize to new tasks in real-time. Particularly, CAMeLU outperforms other UML baselines across all datasets, establishing itself as the state-of-the-art in the field. It also achieves comparable results to its supervised counterpart and to SSL approaches. While the latter requires fine-tuning on the test domain, CAMeLU obtains comparable performance with a single forward step, highlighting its applicability to real-time applications. Furthermore, by recasting the meta-learning phase as in-context learning within a transformer architecture, we improve efficiency, ensuring the whole training and inference phase can be executed with a consumer device with 4GB VRAM.
The main contributions of this paper are as follows:
  • We introduce CAMeLU, a novel UML method that leverages in-context learning within a transformer architecture, reframing meta-learning as a sequence modeling problem.
  • We propose a novel task creation mechanism that generates diverse few-shot tasks from unlabeled datasets using a combination of data augmentations and a mixing strategy. This ensures better alignment between training and testing tasks, thus improving generalization performance.
  • We demonstrate that CAMeLU outperforms existing UML baselines across four datasets, without the need for fine-tuning to the test domains.
  • We investigate the ability of CAMeLU to generalize across various datasets, including those significantly different from the training data.
2 Related work
Unsupervised meta-learning.
Meta-learning is a well-studied field in the machine learning community due to its ability to enable models to quickly adapt to tasks with limited labeled data. Pioneering work in the field [14, 40, 49, 32, 41] considers the scenarios where a large labeled dataset is available for meta-training, a challenging requirement in real-world applications. UML addresses this challenge by extracting meaningful information from unsupervised data that can be transferred to downstream tasks with limited labeled data. Different techniques have been explored in the literature to construct diverse tasks. CACTUs [18] applies clustering in the embedding space and assigns the same pseudo-label to all images in the same cluster. Other methods focus on generating synthetic samples, either using data augmentations, as in UMTRA [20], or leveraging interpolation in the latent space of a generative model [21]. Differently, Meta-GMVAE [25] and Meta-SVEBM [23] use variational autoencoders and memory-based models for pseudo-label generation. In recent methodologies, also SSL approaches [10] have been incorporated into UML methods. In particular, Set-SimCLR [24] builds on top of the SimCLR [7] approach and reframes meta-learning as a set-level problem, while PsCo [19], inspired by MoCo [16], utilizes a momentum encoder and a queue of previous samples to improve pseudo-labeling and construct diverse tasks for UML applications.
Data augmentation.
Several UML approaches rely on data augmentation to construct the training tasks [20, 24, 19]. However, traditional transformations such as rotation, translation, cropping, resizing, and flipping [39] might generate images that are too similar to the original ones, ending up in tasks with low in-class variability between the support and query images. This creates a problem when the model needs to generalize to test tasks, where the query data are different instances than the support ones, not only augmented versions of them. In this paper, we addressed this limitation by generating query images using a strategy inspired by mixup [57] to enhance model generalization. Similarly to mixup [57], which performs linear interpolation of the feature vectors at the pixel level, other strategies based on mixing images comprise CutMix [56], PatchMix [28], and Manifold Mixup [47].
In-context learning.
In-context learning refers to the ability to perform a new task via inference alone by conditioning on a few input-output pairs and making predictions for new inputs [11]. While typical of LLMs [9, 35, 42], this ability has also been explored in different fields, such as in-painting [1, 58], image segmentation [5], and notably meta-learning [22, 13, 12, 31]. In particular, GPICL [22] shows that transformers can be meta-trained to act as general-purpose in-context learners. Similarly, CAML [12] proposes a universal meta-learner by rephrasing the meta-learning framework as a sequence modeling problem and employing a transformer model for few-shot classification. However, this approach is designed for supervised meta-learning, thus requiring labels both during the training and inference phases. Inspired by these, we propose a new approach that integrates the in-context learning ability of transformers with a novel task creation mechanism to solve the UML problem.
3 Proposed approach
Our proposed approach, Context-Aware Meta-Learning in Unsupervised scenarios (CAMeLU), leverages the in-context learning ability of transformers to address the challenges of UML. These challenges include the need to construct meaningful tasks from unlabeled data and the requirement for models to generalize effectively to new tasks during inference. CAMeLU consists of two phases that are intertwined during the model training. Initially, tasks are automatically constructed from an unlabeled dataset utilizing a combination of two strategies. Subsequently, we reformulate the meta-learning framework as a sequence modeling problem, aiming to harness the in-context learning capability of a transformer. This enables the model to extract context from the support samples and predict the unknown query samples without requiring any fine-tuning during the inference phase. We delve into the two phases in Sect. 3.1 and Sect. 3.2, respectively.
3.1 Task creation
Central to our proposed approach is the task creation mechanism. In meta-learning, a task $\mathcal{T}_{i}$ corresponds to a data generating distribution $\mathcal{T}_{i}\triangleq\{p_{i}(x),p_{i}(y|x)\}$, and consists of data from $N$ distinct classes. The data sampled from each task is divided into a support set, $D_{i}^{(sp)}$, containing $K$ training examples per class, and a query set, $D_{i}^{(qr)}$. At meta-test time, only the support set $D_{new}^{(sp)}$ of a task $\mathcal{T}_{new}\sim\mathcal{D}_{test}$ is labeled and used to fine-tune the model and make accurate predictions on the unlabeled query set. Contrary to supervised meta-learning, tasks in UML are only available at test time, while a large unlabeled dataset $\mathcal{D}_{train}$ is available during training. The main goal is to extract prior knowledge from this unlabeled dataset that can be generalized to a target task, $\mathcal{T}_{new}\sim\mathcal{D}_{test}$, during inference. A critical aspect of UML approaches lies in the task creation mechanism to create tasks from $\mathcal{D}_{train}$, which must ensure that the constructed training tasks reflect the structure of those encountered during testing, thereby facilitating effective generalization to novel tasks at test time. To do so, we employ two distinct strategies for constructing the support and query sets of each task.
For the support set, we randomly sample $N$ images from $\mathcal{D}_{train}$ under the assumption that they belong to distinct categories, as shown in Figure 1. This assumption is reasonable when $N<<C$, where $C$ denotes the total number of classes in $\mathcal{D}_{train}$, which is satisfied using a large training dataset. If we assume that all samples are equally distributed among the classes, i.e., $m$ samples per class, the probability that two or more samples are in the same class is equal to
$\textnormal{P}=1-\frac{(C\cdot m)\cdot((C-1)\cdot m)\cdots((C-N+1)\cdot m)}{(C\cdot m)\cdot(C\cdot m-1)\cdots(C\cdot m-N+1)}=\frac{C!\cdot m^{N}\cdot(C\cdot m-N)!}{(C-N)!\cdot(C\cdot m)!}.$
For example, the probability for a 5-way classification on the ImageNet-964 dataset used in our experiment is around $0.01$, which is negligible. To emulate the $K$-shot scenario typical of meta-learning tasks, we augment each of the $N$ images $K$ times, with an augmentation function $\mathcal{A}_{k}$ sampled from a predefined set of transformations $\mathbfcal{A}$, and we assign the same pseudo-label $n\in[1,N]$ to all data generated from the same sample $x_{n}$. Specifically, for each image $x_{n}$, $K$ augmentation functions are applied to obtain $x_{n,k}=\mathcal{A}_{k}(x_{n})$ with $\mathcal{A}_{k}\sim\mathbfcal{A}$ and $k=1,\ldots,K$. One requirement of $\mathcal{A}_{k}$ is that the function must preserve class membership, i.e., $x_{n}\in c\rightarrow\mathcal{A}_{k}(x_{n})\in c$, for $c\in C$. Although this property cannot be directly verified due to the lack of class information in the training set, it is reasonable to assume that it holds by selecting transformations that minimally alter the image content.
For the query set, we employ a different approach. We demonstrate in Appendix 7.2 that simply applying data augmentations sampled from $\mathbfcal{A}$ is not sufficient for creating a query set resembling those in test tasks. At test time, the query set samples are different instances belonging to the same $N$ classes encountered in the support set, not augmented versions of the support samples. However, since $\mathcal{D}_{train}$ is unlabeled, we need a strategy to create new samples with the same implicit classes as those in the support set. We propose a new strategy inspired by mixup [57]. For each query image $x_{j}^{(qr)}$ that we want to generate, we randomly sample an image $z_{j}$ from $\mathcal{D}_{train}$ and another image $x_{i}^{(sp)}$ from the support set. These images are then combined pixel-wise according to:
$x_{j}^{(qr)}=\lambda z_{j}+(1-\lambda)x_{i}^{(sp)}$ (1)
where $\lambda\sim Beta(\alpha,\beta)$ with $\alpha=1,\beta=1$ and $\lambda\in(0,0.5)$, and $x_{j}^{(qr)}$ is assigned the same label as $x_{i}^{(sp)}$. By merging a small proportion of a new image $z_{j}$ into $x_{i}^{(sp)}$, we enhance diversity in the query set with respect to the images in the support set. This strategy enforces the model to extract robust features and effectively generalize to scenarios where query images differ from the support samples, as commonly encountered at test time.
This task-creation mechanism can be seen as a task augmentation strategy [55, 36] that allows the generation of a large number (almost infinite) of diverse tasks. This is particularly useful for meta-learning and in-context learning applications where the model needs to acquire knowledge from a multitude of tasks to generalize to unseen tasks sampled from different domains.
Differences with mixup.
While the strategy used for generating the query images draws inspiration from the mixup strategy proposed in [57], there are some substantial differences. The aim of mixup [57] is to develop a new data augmentation strategy to expand the number of training examples and diversify the data distribution used for training, thereby enhancing the robustness and generalization of neural networks. In CAMeLU, the primary objective of merging images is to encourage the model to learn even in scenarios where only a fraction of the class context is present in the image. In CAMeLU, $\lambda$ is sampled from a uniform distribution (obtained with a $Beta$ distribution with $\alpha=1,\beta=1$) in $(0,0.5)$, guaranteeing that the amount of information from $z_{j}$ that is embedded into $x_{j}^{(qr)}$ is less than 50%, which would otherwise corrupt the label data. Indeed, we assign the same label of $x_{i}^{(sp)}$ to $x_{j}^{(qr)}$, forcing the network to learn to retrieve information in $x_{j}^{(qr)}$ that is related to the category of $x_{i}^{(sp)}$. Contrarily, mixup creates new examples by interpolating both images and labels at the scope of extending the training distribution.
3.2 In-context learning method
Following task creation, we rephrase the meta-learning framework as a sequence modeling problem. Inspired by recent developments in LLMs [15, 26, 9, 35, 42], we treat each task as a prompt, where the support embedding, together with the learned projected labels, form the demonstration context, whereas the query represents the classification problem that the network is required to solve. A model is said to in-context learn a task if it can approximate $y_{j}^{(qr)}$ for a new query input $x_{j}^{(qr)}$ by conditioning on a sequence $S_{i,j}$ containing in-context (support) examples and one query input defined as follows:
$S_{i,j}=\left((x_{1}^{(sp)},y_{1}^{(sp)}),\ldots,(x_{NK}^{(sp)},y_{NK}^{(sp)}),x_{j}^{(qr)}\right),\quad j=1,\ldots,Q,$ (2)
with $Q$ the number of query samples to classify and $NK$ the total number of context (support) samples. Formally, $M_{\theta}$ can in-context learn a task $\mathcal{T}_{i}$ if it can predict $y_{j}^{(qr)}$ with an average error
$\mathbb{E}\left[\sum_{j=1}^{Q}\ell(M_{\theta}(S_{i,j}),y_{j}^{(qr)})\right]<\epsilon,$ (3)
where $\ell$ is the loss function, $S_{i,j}$ is the sequence associated to $x_{j}^{(qr)}$ in $\mathcal{T}_{i}$, and $y_{j}^{(qr)}\in[1,N]$.
To achieve this, we design a model comprising three components: (1) a feature extractor $f_{\psi}$, (2) a class encoder $g_{\phi}$, and (3) a transformer encoder with a linear projection layer on top, i.e., $M_{\theta}$. The feature extractor aims to map support and query samples into a latent space where images with similar characteristics and semantic meaning are assigned similar representations. Section 7.4 explores various feature extractors for this purpose, including those pre-trained via a supervised approach or leveraging an SSL technique. The resulting representations are then concatenated with a class embedding. The class embeddings for the support representations are generated by encoding the corresponding classes using the class encoder $g_{\phi}$. However, as the classes of the queries are unknown, a randomly initialized learnable vector is appended to each query representation. The so-combined embeddings are then organized into sequences $S_{i,j}=\left((f_{\psi}(x_{1}^{(sp)}),g_{\phi}(y_{1}^{(sp)})),\ldots,(f_{\psi}(x_{NK}^{(sp)}),g_{\phi}(y_{NK}^{(sp)})),f_{\psi}(x_{j}^{(qr)})\right),$ $j=1,\ldots,Q,$ resembling the one in Equation 2. These sequences are fed into the transformer encoder, and only the transformer output corresponding to the query sample is selected and passed through a projection layer to predict the query label. This process iterates for all queries in the task, and the aggregated loss is employed for model training. In particular, the training process can be formulated as an optimization process where the objective is as follows:
$\min_{\theta,\phi}\;\mathbb{E}_{S_{i}}\left[\frac{1}{Q}\sum_{j=1}^{Q}\ell(M_{\theta}(S_{i,j}),y_{j}^{(qr)})\right]$ (4)
with $S_{i}=\{S_{i,j}\}_{j=1}^{Q}$ denoting the set of sequences associated to each task $\mathcal{T}_{i}$ generated from $\mathcal{D}_{train}$ and $\ell$ is the cross-entropy loss function.
4 Experiments
In this section, we demonstrate the effectiveness of CAMeLU across different datasets, and we compare the results with several baseline methods. In particular, we provide a quantitative comparison with UML baselines in Sect. 4.3, and we highlight the ability of CAMeLU to leverage generalization over memorization in Sect. 4.4. We then present the results with a large-scale training dataset in Sect. 4.5 and a comparison with SSL methods in Sect. 4.6.
4.1 Datasets and baselines
For the evaluation, we use two generic object recognition datasets, i.e., miniImageNet [37] and CIFAR-fs [2], and two fine-grained image classification datasets, i.e., CUB [50], and Aircraft [30]. Each dataset is split into training, validation, and test sets following the splits in [37] and [2] for miniImageNet and CIFAR-fs, respectively, and in [43] for the remaining datasets. All labels are removed from the datasets during the training phase.
We compare the results of CAMeLU with standard UML approaches such as CACTUs [18], UMTRA [20], Meta-GMVAE [25], and PsCo [19]. These methods are evaluated in-domain as recommended in the original papers, with training and testing performed on the same dataset. While this setup is relatively simpler than the cross-domain evaluation employed for CAMeLU, applying these methods in a cross-domain scenario may not be fair, as they were not explicitly designed for such a challenging scenario. Only PsCo [19] is further evaluated in a cross-domain setting, as the authors demonstrate its adaptability to this scenario through an additional adaptation phase to the test domain. The performance of CAMeLU is also compared with CAML [12], a supervised meta-learning approach that assumes tasks are available both during the training and testing phases and leverages the in-context ability of transformer architectures to generalize to new tasks. Due to their similarities, we refer to CAML as the supervised counterpart of CAMeLU. Furthermore, Section 4.6 reports a comparative analysis with two fine-tuned state-of-the-art self-supervised trained networks, namely SimCLR [7] and SwAV [6].
4.2 Training details
We report the results following the $N$-way $K$-shot classification task typical of meta-learning algorithms, where $N=5$ and $K=1$ or $K=5$. All the models are trained for $100$ epochs, $500$ episodes per epoch, and $100$ fine-tuning steps at test time if required by the method. For CAMeLU, we choose not to apply any fine-tuning strategy to prove the strength of the training stage that does not require additional parameter updates during the inference phase. To validate the cross-domain performance, we trained the models on ImageNet-964 after removing the image labels. This is a variant of the original ImageNet-1k [8] where classes belonging to the validation and test splits of miniImageNet are removed. This is done to prevent any data leakage when evaluating the model on miniImageNet - a problem that is not taken into consideration in previous studies [12, 19]. For CAMeLU, we use a feature extractor pre-trained using CLIP [34] on a ViT-B/16 architecture and a class encoder consisting of a single learnable layer. In Appendix 7.4, we also report the results with different feature extractors. The transformer encoder consists of 8 encoder layers, each incorporating a multi-head self-attention block with 8 attention heads, an MLP with a reverse bottleneck of $3072$ and an input-output feature size of $1024$, where the last $768$ features are the ViT feature vector appended to the $256$ encoding of the labels. Finally, a single projection layer maps the transformer output to the predicted category. These choices were motivated by our hardware limitations, which constrained us to load and train the model on a consumer device with only 4GB of VRAM. The model is trained with the Adam optimizer with a learning rate of $10^{-5}$ and a warmup cosine scheduler as proposed in the original paper [46]. To account for statistical variations, each algorithm is run three times in full. The experiments are executed using Python and the PyTorch library on an Nvidia GeForce RTX 3050 Ti Laptop GPU with 4GB of VRAM. Further ablation experiments and competitors required a more capable architecture and are executed on an Nvidia A100-SXM4 GPU with 40GB of VRAM. More details about the training settings can be found in Appendix 7.1, while the code will be released upon acceptance of the paper.
4.3 Comparative results
Table 1 provides an overview of the experimental results for both the 5-way 1-shot and the 5-way 5-shot scenarios. The results demonstrate that CAMeLU outperforms the existing UML methods, regardless of the difference in the evaluation setting. As highlighted in Section 4.1, CACTUs, UMTRA, and Meta-GMVAE are evaluated only in-domain, requiring knowledge about the test domain prior to training. This is not necessary for CAMeLU as it demonstrates high performance in the challenging cross-domain scenario. Even compared to PsCo, the only UML method designed for cross-domain applications, CAMeLU exhibits a performance improvement in all datasets except miniImageNet, where PsCo achieves even better results than the supervised CAML. This is expected, as the model utilized for cross-domain evaluation of PsCo is pre-trained on ImageNet-1k, which includes the images of miniImageNet, leading to data leakage. Furthermore, PsCo requires a fine-tuning phase to adapt to the test domain, whereas CAMeLU achieves good performance with a single forward pass, enhancing its applicability to real-time applications. It is also worth noting that CAMeLU achieves comparable performance to its supervised counterpart, CAML, when evaluated on miniImageNet and CIFAR-fs and it even outperforms CAML when evaluated on more dissimilar domains, such as CUB and Aircraft. This highlights the efficacy of the task construction strategy used in CAMeLU, which acts as a sort of task augmentation and enhances the generalization capability of the model.
Table 1: Performance comparison on miniImageNet, CIFAR-fs, CUB, and Aircraft dataset for 5-way 1-shot and 5-way 5-shot scenarios. In-domain approaches are trained with a Conv5 architecture, while PsCo cross-domain is trained on a ResNet-50 architecture. The symbol $\dagger$ indicates results that are affected by data leakage. The bold font highlights the best performing UML approach for each setting. Results show the mean and standard deviations across three complete runs of the algorithms.
miniImageNet CIFAR-fs CUB Aircraft Method 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s In-Domain CACTUs-MAML $43.30\pm 0.29$ $54.21\pm 1.00$ $42.00\pm 1.47$ $56.64\pm 0.49$ $31.19\pm 0.37$ $36.81\pm 0.68$ $24.06\pm 0.78$ $27.26\pm 0.04$ CACTUs-ProtoNet $48.85\pm 0.69$ $62.52\pm 0.71$ $50.90\pm 0.46$ $64.52\pm 0.94$ $33.93\pm 0.37$ $44.41\pm 1.31$ $26.27\pm 0.28$ $30.88\pm 0.51$ UMTRA $39.93\pm 1.15$ $50.73\pm 0.67$ $32.93\pm 1.68$ $46.13\pm 2.81$ $27.06\pm 1.41$ $36.6\pm 2.43$ $22.40\pm 3.42$ $31.73\pm 2.25$ Meta-GMVAE $55.38\pm 0.90^{\dagger}$ $65.10\pm 0.64^{\dagger}$ $52.02\pm 0.88$ $64.18\pm 0.62$ $33.59\pm 0.63$ $39.09\pm 0.57$ $24.83\pm 0.51$ $27.60\pm 0.52$ PsCo $45.80\pm 0.01$ $62.64\pm 0.01$ $51.62\pm 0.01$ $68.43\pm 0.01$ $35.90\pm 0.01$ $47.46\pm 0.01$ $26.10\pm 0.01$ $31.40\pm 0.01$ Cross-Domain PsCo $\mathbf{82.08\pm 0.42}^{\dagger}$ $\mathbf{94.61\pm 0.15}^{\dagger}$ $\mathbf{66.28\pm 0.50}$ $83.36\pm 0.34$ $52.16\pm 0.53$ $75.35\pm 0.44$ $27.17\pm 0.32$ $42.01\pm 0.43$ CAMeLU $69.63\pm 2.87$ $90.95\pm 0.54$ $58.60\pm 1.92$ $\mathbf{84.26\pm 0.77}$ $\mathbf{55.12\pm 1.42}$ $\mathbf{77.71\pm 3.17}$ $\mathbf{36.93\pm 0.53}$ $\mathbf{46.27\pm 2.18}$ CAML (supervised) $92.58\pm 0.12$ $97.93\pm 0.05$ $66.33\pm 3.22$ $83.39\pm 1.28$ $45.76\pm 0.81$ $68.40\pm 2.55$ $33.04\pm 1.82$ $35.82\pm 0.92$
4.4 Memorization to generalization phase shift
During the training of CAMeLU, we observed a distinct trend in the validation accuracy, similar to the findings in [22]. Figure 2 illustrates this pattern, showing the validation accuracy relative to its initial value, or, in other words, how much the model learns from datasets different from the one we are training on. Specifically, the curves in Figure 2 resemble a logistic curve, which can be divided into three phases that we denote as memorization, learning, and generalization. In the memorization phase, the model memorizes the tasks seen during training and extends this knowledge to unseen tasks, resulting in a slight improvement for datasets with high similarity with ImageNet-964 (e.g., miniImageNet). For the other datasets, instead, transferring this knowledge can even result in a performance decrease due to the intrinsic domain distance of the dataset (see CUB and Aircraft, which are fine-grained datasets). As training progresses and the model observes more tasks, the learning phase occurs. This phase is characterized by a transition to the learning-to-learn state where the model learns to identify the tasks and to extract the features that are more useful for solving them.
Figure 2: Analysis of learning behavior when transferring knowledge from different prior datasets. The relative validation accuracy shows the difference between the current and first epoch accuracy on the validation set of miniImageNet, CIFAR-fs, CUB, and Aircraft. CAMeLU is trained with ImageNet-964.
The duration of this phase varies, with datasets like miniImageNet and CIFAR-fs exhibiting rapid learning within approximately $10$ epochs, while datasets such as CUB and Aircraft may necessitate up to 40 epochs. This timespan depends on several factors, including the similarity between the training and evaluation datasets, the size of the test dataset, and the model’s learning ability [22, 33]. For instance, CUB, with its fine-grained nature and small test set size (around $1770$ images) necessitates a longer learning phase compared to the miniImageNet dataset (which has a test set with 12 0001200012\,00012 000 images). Subsequently, in the generalization phase, the model can generalize to tasks significantly different from those observed during training using a single forward pass. Further analyses about the generalization capabilities of CAMeLU and the number of epochs required for reaching the generalization phase are presented in Appendices 7.5 and 7.6.
4.5 Multi-dataset training
Table 2: Comparison of CAMeLU and CAML when trained on the single ImageNet-964 dataset (denoted as IN-964) and on a multi-dataset consisting of ImageNet-964 + MSCOCO + Fungi. Results show the mean and standard deviations for the 5-way 1-shot and the 5-way 5-shot settings across three complete runs of the algorithms.
miniImageNet CIFAR-fs CUB Aircraft 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s CAMeLU (IN-964) $69.63\pm 2.87$ $90.95\pm 0.54$ $58.60\pm 1.92$ $84.26\pm 0.77$ $55.12\pm 1.42$ $77.71\pm 3.17$ $36.93\pm 0.53$ $46.27\pm 2.18$ CAMeLU (mds) $70.42\pm 2.77$ $91.02\pm 0.97$ $60.86\pm 3.94$ $\mathbf{85.05\pm 0.27}$ $\mathbf{55.58\pm 2.68}$ $\mathbf{78.68\pm 2.13}$ $\mathbf{38.88\pm 2.22}$ $\mathbf{48.86\pm 3.45}$ CAML (IN-964) $\mathbf{92.58\pm 0.12}$ $\mathbf{97.93\pm 0.05}$ $66.33\pm 3.22$ $83.39\pm 1.28$ $45.76\pm 0.81$ $68.40\pm 2.55$ $33.04\pm 1.82$ $35.82\pm 0.92$ CAML (mds) $92.10\pm 0.33$ $97.82\pm 0.03$ $\mathbf{67.01\pm 2.66}$ $\mathbf{85.05\pm 2.88}$ $44.03\pm 2.93$ $73.47\pm 2.56$ $31.51\pm 0.58$ $38.88\pm 4.54$
We conduct additional experiments to evaluate how CAMeLU performs when trained on a large-scale dataset. As our focus is cross-domain classification through in-context learning, we hypothesize that training on a dataset spanning various concepts enhances classification performance, as suggested in [12, 31]. To this effect, we combine three training datasets with different levels of granularity: ImageNet-964 [8], MSCOCO [27], and Fungi [38]. At each training episode, a dataset is uniformly sampled, and $N$ data points are subsequently extracted. These samples are then utilized in our method as described in Section 3. Table 2 presents the results for CAMeLU and the supervised CAML method. Notably, training on a combination of multiple datasets (denoted as mds in Table 2) yields improved performance compared to training solely on ImageNet-964. However, the performance gap is not substantial, likely due to the large size of ImageNet-964 (around $80\%$ of the overall dataset), resulting in its more frequent selection compared to the other datasets. Therefore, to optimize computational resources and time, we conducted all the remaining experiments by training only on ImageNet-964.
4.6 Comparison with SSL methods
In this section, we compare CAMeLU with SimCLR [7] and SwAV [6]. For training SSL methods in our experiments, we employed a backbone network with a ResNet-50 architecture pre-trained on ImageNet-1k and obtained from PyTorch Lightning Bolts [3]. To facilitate model adaptation to the test domain, we fine-tuned a classification layer on top of the pre-trained backbone using SGD with initial learning rate of $0.01$, momentum of $0.9$, weight decay of $10^{-4}$, and $100$ fine-tuning steps per task, following [19]. This setup differs from the evaluation setting used in CAMeLU, where predictions are obtained with a single forward pass, leveraging the in-context learning ability of transformer architectures. However, SSL approaches must adapt to the test domain before label predictions, resulting in a less challenging evaluation setting than CAMeLU. Results for SSL approaches are averaged over $500$ test tasks and presented in Table 3. While SSL approaches outperform CAMeLU on miniImageNet and CIFAR-fs, their performance decreases when evaluated on the other datasets. CUB and Aircraft are two fine-grained datasets significantly different from ImageNet-1k, challenging the transferability of features learned by SSL methods to these datasets. Moreover, the high performance on miniImageNet may be attributed to data leakage with ImageNet-1k, as discussed in Section 4.2. CAMeLU, in contrast, demonstrates effective generalization to tasks sampled from these datasets, once again highlighting its generalization ability over mere memorization.
Table 3: Comparison between CAMeLU and SSL approaches for the 5-way 1-shot and 5-way 5-shot scenario on miniImageNet, CIAFR-fs, CUB, and Aircraft. The symbol $\dagger$ indicates results that are affected by data leakage. Results show the mean and standard deviations across three complete runs of the algorithms.
miniImageNet CIFAR-fs CUB Aircraft Method 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s SimCLR $\mathbf{83.32\pm 0.23}^{\dagger}$ $\mathbf{94.86\pm 0.61}^{\dagger}$ $\mathbf{64.52\pm 0.69}$ $84.36\pm 0.40$ $47.35\pm 0.53$ $66.87\pm 0.82$ $29.36\pm 0.90$ $39.99\pm 0.86$ SwAV $74.83\pm 0.71^{\dagger}$ $\mathbf{94.96\pm 0.91}^{\dagger}$ $66.97\pm 0.15$ $\mathbf{87.14\pm 0.10}$ $47.84\pm 0.31$ $69.31\pm 0.01$ $30.33\pm 0.31$ $\mathbf{47.43\pm 0.11}$ CAMeLU $69.63\pm 2.87$ $90.95\pm 0.54$ $58.60\pm 1.92$ $84.26\pm 0.77$ $\mathbf{55.12\pm 1.42}$ $\mathbf{77.71\pm 3.17}$ $\mathbf{36.93\pm 0.53}$ $46.27\pm 2.18$
5 Limitations
Despite the promising results demonstrated by our novel UML approach on several datasets, some limitations remain. Its applicability and robustness in real-world scenarios with diverse and noisy data remain to be thoroughly evaluated. In real-world applications, data can be incomplete, mislabeled, or drawn from significantly different distributions, leading to potential degradation in the model’s performance in the presence of noisy or corrupted data. Additionally, CAMeLU’s performance depends on the pre-trained feature extractor, as demonstrated in Appendix 7.4. Replacing the extractor pre-trained with CLIP with one pre-trained using an SSL strategy could make the pipeline completely unsupervised, albeit at the price of performance degradation. Lastly, the proposed approach is designed to handle a fixed number of classes (ways) per task during training and testing, requiring knowing the value of $N$ in advance. Modifications to the task creation and training process would be necessary to extend our approach to handle an arbitrary number of ways.
6 Conclusion
In this paper, we introduce CAMeLU, a novel approach to UML that leverages the in-context learning capabilities of transformer architectures to extract context from the support samples and make effective predictions on the query data. CAMeLU reframes meta-learning as a sequence modeling problem, where support images provide task context for predicting query images. At the core of CAMeLU is a novel task creation mechanism that generates diverse tasks from an unlabeled dataset, promoting effective generalization to unseen tasks. Our experimental results showcase the superiority of CAMeLU over existing UML methods, highlighting the applicability of the proposed method to domains different from the training one. Notably, CAMeLU can generalize to new domains with a single forward pass (real-time predictions), and it even outperforms its supervised counterpart thanks to its task creation mechanism. Furthermore, the proposed model can be stored and trained with a single GPU with only 4GB of VRAM, underscoring its efficiency in learning-to-learn in-context, rather than using a meta-training phase typical of previous meta-learning approaches.
Future research directions may explore extensions of CAMeLU to more complex domains, as well as investigations into further improving the task creation mechanism for enhanced generalization. It would be interesting to incorporate SSL techniques to obtain more robust feature representations and enhance generalization capabilities. Additionally, conducting further investigation into CAMeLU’s ability to encourage generalization over memorization would provide valuable insights into its learning dynamics and potential areas for improvement.